There are 150 students in a class. The distribution if their marks in a mathematics test are as follows

Class                                                   frequency

0-9                                                      3

10-19                                                  10

20-29                                                  17

30-39                                                  x

40-49                                                  35

50-59                                                  y

60-69                                                  18

70-79                                                  10

80-89                                                  5

90-99                                                  2

Required

i)       The value of x given that the median mark is 44.357            (2marks)

ii)     The value of y given that the modal is 43.0                          (2marks)

iii)   Draw an ogive of the data in (a) above                                 (3 marks)


1
Expert's answer
2022-03-17T06:34:21-0400



i. median "=44.357=L+(\frac{(\frac{N}{2}-pcf)}{f})i"

"=40+(\frac{(75-(30-x)}{35})10"

"=39.5+(\frac{(45-x)}{35})10"

"4.357=(\frac{(45-x)}{35})10"

"0.4357=\frac{(45-x)}{35}"

"15.2495=45-x"

"x=29.75"

ii. Mode "=43.0=40+(\frac{7}{(7+35-y)})10"

"40+(\frac{7}{(42-y)})10"

"3=(\frac{7}{(42-y)})10"

"0.3=\frac{7}{(42-y)}"

"42-y=\frac{7}{0.3}=23.33"

"y=42-2.33=18.67"

iii.